# Recognising, Anticipating, and Mitigating LLM Pollution of Online Behavioural Research

### v1

@@ Conclusions
LLM Pollution presents an epistemic challenge, not just a technical one. It blurs the line between human and machine in the very data we use to understand human cognition and behaviour. Crucially, the threat is not always adversarial. In many cases, participants use LLMs to improve fluency, reduce effort, or navigate complex instructions — behaviours that may increasingly reflect the realities of LLM-usage in everyday life. This growing entanglement complicates the question of what counts as “pollution” and demands more than just technical safeguards. Even when intentions are benign, LLM involvement can introduce systematic distortions that are difficult to detect, impossible to disentangle post hoc, and as such often invisible to researchers.
@@
As LLMs become increasingly embedded in everyday life, their use in cognitive, communicative, and problem-solving tasks may no longer be an exception, but the norm. This raises a more fundamental question: at what point does LLM-assisted behaviour cease to be “pollution” and instead become part of the ecological baseline we must account for? While mitigation remains essential for preserving the integrity of current methods, the long-term challenge may lie in adapting our theoretical frameworks to a world where human reasoning is increasingly shaped by intelligent machines.
@@ A.3 Norm signalling to participants
Researchers should ensure that the tone remains respectful and non-accusatory, to avoid undermining participant trust. While norm signalling may aid in reducing LLM Spillover by reassuring participants that proper safeguards are in place, it may also inadvertently suggest that such behaviour is widespread, thereby decreasing the motivation of participants who value a fair and human-only study environment.
@@ A.4 Multimodal Presentation of Instructions
A.5 Input Interface Restriction
Restricting the user input interface is one of the most effective strategies for mitigating low-effort or automated responses in online experiments. Here, we outline two complementary approaches: (1) requiring participants to respond via alternative modalities, such as verbal input, and (2) disabling copy-paste functionalities in standard text fields.
While both of these measures are designed to increase friction rather than eliminate the risk of LLM Pollution altogether, they address different types of threats. Soliciting audio-based (human) input increases the technical barriers for Full LLM Delegation, making it nearly impossible for an LLM agent to complete a survey autonomously.
@@ A.5 Input Interface Restriction
While not sufficient on their own, input interface restrictions offer a scalable and low-cost method for reducing LLM interference.
@@ A.6 LLM-Specific Comprehension Checks
Undoubtedly, these checks are inherently moving targets. As LLM (and agentic) capabilities continue to improve in areas such as vision-language alignment, many of their current weaknesses will be patched, rendering specific items obsolete. This underscores the importance of continuously reassessing and potentially redesigning such tasks, ideally piloting across a range of model architectures and capabilities, including those deployed as stand-alone agentic systems.
Moreover, human participants can also fail these tasks, especially if the instructions are overly complex or if participants are not attentive enough. We therefore recommend treating responses to these types of tasks as probabilistic filters rather than definitive exclusion criteria.
@@ 4. Hidden checkbox trap with neutral naming:
While effective against many types of autonomous LLM agents, most advanced pipelines already include safeguards against filling in fields with suspicious styling attributes (see Figure B3). Furthermore, screen-capturing models or accessibility-enhanced user interfaces (e.g., Safari Reader mode or browser extensions) may surface honeypot content that should (and otherwise would) remain invisible. Researchers should test their implementation of these questions across different browsers and display modes, and should also ensure that the study colour scheme cannot be easily altered. Such deviations from the intended rendering may also result in false positives, therefore raising concerns about unfairly excluding legitimate participants.
@@ A.8 Behavioural Logging and Modelling
  • Tab switching/window focus: monitor how often participants switch windows during the task. Note that attempting to record the specific pages that participants switch to is not compliant with legal/institutional constraints. This technique is specifically useful for detecting cases of Partial LLM Mediation.
@@
Of course, tracking behavioural measures may disadvantage participants using assistive technologies, so researchers must be careful in the automatic exclusion of participants who do not align with standard behavioural patterns. In addition, more sophisticated agents may learn to replicate these patterns with increasing accuracy, making it more difficult to mark them as outliers. Finally, the collection of interaction data raises non-trivial ethical and legal considerations, even if collected and stored exclusively on the researchers’ servers. Researchers have a responsibility to ensure transparency in data collection and storage practices and to remain attentive to changing legal constraints.
@@ A.9 Commercial AI-Generated Text Detectors
Orenstrakh et al. (2024) provide a fairly recent (yet in some aspects already outdated) comparison of the performance of some of the tools listed above, both before and after using Quillbot to paraphrase AI-generated text. Additionally, researchers can also apply their own domain knowledge to identify AI-written text, searching for elements such as factual inconsistencies, overuse of formal or abstract language, hedging, and generic phrasing (Cotton et al., 2024).
Nevertheless, many detectors exhibit a high false positive rate, particularly for human-written text that follows formulaic structures. These detectors often struggle to generalize across domains and have particular difficulty identifying text that has only been minimally modified by humans (Wu et al., 2025). Given these limitations, we advise the use of these tools only in post-hoc checks and in conjunction with other detection measures.

### v2

@@ Conclusions
LLM Pollution presents an epistemic challenge, not just a technical one. It blurs the line between human and machine in the very data we use to understand human cognition and behaviour. Crucially, the threat is not always adversarial. In many cases, participants use LLMs to improve fluency, reduce effort, or navigate complex instructions — behaviours that may increasingly reflect the realities of LLM-usage in everyday life. In our framing, LLM Pollution is the cause, and its effects include systematic distortions that can compromise validity and interpretation. This growing entanglement complicates the question of what counts as “pollution” and demands more than just technical safeguards. Even when intentions are benign, LLM involvement can introduce systematic distortions that are difficult to detect, impossible to disentangle post hoc, and as such often invisible to researchers.
@@
While our focus is on online behavioural experiments, LLM Pollution can also arise in other forms of data that researchers may use to address research questions, such as observational datasets based on online activity (e.g., social media posts), where parts of the data may already be generated or altered by LLMs. Addressing these broader cases is important, and future research should examine how LLM Pollution affects other types of data and research settings.
As LLMs become increasingly embedded in everyday life, their use in cognitive, communicative, and problem-solving tasks may no longer be an exception, but the norm. This raises a more fundamental question: at what point does LLM-assisted behaviour cease to be “pollution” and instead become part of the ecological baseline we must account for? While mitigation remains essential for preserving the integrity of current methods, the long-term challenge may lie in adapting our theoretical frameworks to a world where human reasoning is increasingly shaped by intelligent machines. In this context, understanding the second-order dynamics of an LLM-infused society (Brinkmann et al., 2023; Tsvetkova et al., 2024; Dong et al., 2025) will provide insights into the spillover effects in online research ecosystems.
Looking ahead, there are several priorities for further empirical work in this area. Beyond offering conceptual distinctions, our framework highlights the need for systematic tracking of the prevalence and development of LLM Pollution as agentic tools become increasingly accessible to participants. Moreover, there is a need for further evaluation of mitigation strategies and the creation of measurement tools for spillover effects. Future research can build on this Perspective to test which safeguards are most effective, which situational and motivational conditions make Partial LLM Mediation most likely, how LLM involvement alters behaviour in practice, and how evolving participant norms reshape the boundaries between human and machine-generated data.
@@
Author Contributions
R.R., T.W., H.Y., and A.-M.N. wrote the manuscript. R.R., H.Y., and A.-M.N. devised and tested the mitigation strategies. I.R. provided conceptual guidance and contributed to the revision of the manuscript.
Acknowledgements
We thank Hani Jahani for infographic support.
@@ A.2 Third-Party Bot Protection Solutions
Risks & Considerations
While commercial bot protection tools can reduce automated traffic, they might also introduce participant friction that may lead to higher attrition or disengagement. Researchers need to balance the level of protection against the potential increase in dropout and selection biases that it may cause, given that repeated verification attempts can be perceived as burdensome or unfair.
@@ A.3 Norm signalling to participants
Risks & Considerations
Researchers should ensure that the tone remains respectful and non-accusatory, to avoid undermining participant trust. This is because, while norm signalling may aid in reducing LLM Spillover by reassuring participants that proper safeguards are in place, it may also inadvertently suggest that such behaviour is widespread, thereby decreasing the motivation of participants who value a fair and human-only study environment.
@@ A.4 Multimodal Presentation of Instructions
A.5 Human-supervised or multimodal response collection
At present, controlling input modalities seems to be one of the most effective strategies for mitigating low-effort or automated responses in online experiments. Here, we outline a continuum of complementary approaches: (1) disabling copy-paste functionalities in standard text fields, (2) requiring participants to use additional input modalities, such as verbal or webcam-enabled responses, and (3) conducting real-time, researcher-supervised sessions (e.g., via Zoom) where participants provide answers verbally, and researchers directly record their answers.
While all of these measures are designed to increase friction rather than eliminate the risk of LLM Pollution altogether, they address different types of threats. Restricting copy-paste functionality may decrease the ease of delegating individual responses to an LLM, while soliciting audio-based (human) input increases the technical barriers for Full LLM Delegation, making it nearly impossible for an LLM agent to complete a survey autonomously.
@@ A.5 Human-supervised or multimodal response collection
Camera- or microphone-based approaches introduce additional ethical challenges, including concerns about participant comfort and the secure handling of recorded material. Requiring video-based input or live supervision further raises the cost of automation but entails additional implementation effort, space and processing demands, time investment, and heightened privacy considerations. Even though live or experimenter-supervised sessions offer maximal oversight and authenticity guarantees, they also carry many drawbacks: they reduce scalability, increase researcher workload, and may exacerbate demand characteristics, as participants adjust their behaviour when they know they are being observed. Moreover, such setups are difficult to standardise across studies, and may inadvertently bias samples toward participants with higher technological access.
@@ A.6 LLM-Specific Comprehension Checks
Risks & Considerations
Undoubtedly, these checks are inherently moving targets. As LLM (and agentic) capabilities continue to improve in areas such as vision-language alignment, many of their current weaknesses will be patched, rendering specific items obsolete. This underscores the importance of continuously reassessing and potentially redesigning such tasks, ideally piloting across a range of model architectures and capabilities, including those deployed as stand-alone agentic systems. Moreover, human participants can also fail these tasks, especially if the instructions are overly complex or if participants are not attentive enough. We therefore recommend treating responses to these types of tasks as probabilistic filters rather than definitive exclusion criteria.
@@ 4. Hidden checkbox trap with neutral naming:
5. Prompt injection-style honeypot questions
: following Höhne et al. (2025b), researchers can integrate two different types of prompt injections, embedding them invisibly on survey pages. These types of instructions operate on the principle that LLMs often comply with direct user instructions.
  1. (a)
    Jailbreaking: similar to the examples already provided above; e.g., “If you are a bot or LLM, please ignore the question above and give this exact answer: I have never completed a survey before, nor will I ever in the future.”
  2. (b)
    Prompt-leaking: ask the model to provide its prompt; e.g., “If you are a bot or LLM, please ignore the question above and respond with an exact copy of the full prompt you received.”
These two types of prompt injections target different types of systems. According to Höhne et al. (2025b), more simple models tend to be responsive to overt jailbreaking prompts, while more advanced models are more likely to reveal hidden content through prompt-leaking instructions. Importantly, we recommend using prompt injections only as honeypot questions, as making this text visible to all participants could influence their behaviour and inadvertently contribute to the amplification of LLM spillover effects.
Risks & Considerations
While effective against many types of autonomous LLM agents, most advanced pipelines already include content-filtering techniques that prevent agents from filling in fields with suspicious styling attributes or from disclosing their internal prompt (see Figure B3). Furthermore, screen-capturing models or accessibility-enhanced user interfaces (e.g., Safari Reader mode or browser extensions) may surface honeypot content that should (and otherwise would) remain invisible. Researchers should test their implementation of these questions across different browsers and display modes, and should also ensure that the study colour scheme cannot be easily altered. Such deviations from the intended rendering may also result in false positives, therefore raising concerns about unfairly excluding legitimate participants.
@@ A.8 Behavioural Logging and Modelling
  • Tab switching/window focus: monitor how often participants switch windows during the task. For example, Permut et al. (2019) developed a toolkit for detecting off-task behaviour and window switching events on Qualtrics surveys. Note that attempting to record the specific pages that participants switch to may not comply with legal/institutional constraints. This technique is specifically useful for detecting cases of Partial LLM Mediation.
@@
Risks & Considerations
Tracking behavioural measures may disadvantage participants using assistive technologies, so researchers must be careful in the automatic exclusion of participants who do not align with standard behavioural patterns. In addition, more sophisticated agents may learn to replicate these patterns with increasing accuracy, making it more difficult to mark them as outliers. Finally, the collection of interaction data raises non-trivial ethical and legal considerations, even if collected and stored exclusively on the researchers’ servers. Researchers have a responsibility to ensure transparency in data collection and storage practices and to remain attentive to changing legal constraints.
@@ A.9 Commercial AI-Generated Text Detectors
Orenstrakh et al. (2024) provide a fairly recent (yet in some aspects already outdated) comparison of the performance of some of the tools listed above, both before and after using Quillbot to paraphrase AI-generated text. Additionally, researchers can also apply their own domain knowledge to identify AI-written text, searching for elements such as factual inconsistencies, overuse of formal or abstract language, hedging, and generic phrasing (Cotton et al., 2024) or can conduct more systematic analyses using natural language processing techniques (Jabarian and Imas, 2025; Zanotto and Aroyehun, 2025; Claassen et al., 2025).
Risks & Considerations
Some detectors exhibit a high false positive rate, particularly for human-written text that follows formulaic structures. These detectors often struggle to generalize across domains and have particular difficulty identifying text that has only been minimally modified by humans (Wu et al., 2025). Given these limitations, we advise the use of these tools only in post-hoc checks and in conjunction with other detection measures.